Diffusion Morning Report 01/04/06

Ted Berg tberg at snf.stanford.edu
Wed Jan 4 06:57:18 PST 2006


Good Morning All,
    Welcome back.  Well we are still trying to recover in Diffusion 
Land. All oxides had TLC clean and should be OK.  LPs had tests but 
results were not exactly what we expected-SiH4 did not flow on poly and 
LTO so they are running again.  Epi is down waiting for burn box 
rebuild. Hopefully we will be back to full speed ahead by a bit later 
this morning.

  	Date: 	1/4/2006 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Up 	  	 
Tylan 8-SiGe 	Up 	  	13.5/30
Tylan 9-Poly 	yellow 	testing 	7.76/30
Tylan 10-Nitride 	yellow 	testing 	13.27/30 
Tylan 11-BPSG-LTO 	yellow 	testing 	15.52/30
Tylan 12-FGA 	Up 	  	 
  	  	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing 	 
Thermco Nitride 	Down 	Testing 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	  
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	A on line 160 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060104/4d2a03fe/attachment.html>


More information about the diffusion mailing list